# BiMark: Unbiased Multilayer Watermarking for Large Language Models

### v1

@@ 5.1 Message Embedding Capacity
Experimental setup. For experiments of message embedding capacity, the Llama3-8B model (AI@Meta, 2024) is used for text generation with temperature $1.0$ and top-$50$ sampling. C4-RealNewslike (Raffel et al., 2020) dataset is used as prompts. For pseudorandom operations of watermarking methods, a 2-token sliding context window is used for seeding. For evaluating the generation quality of language models with watermarks, perplexity of generated text is calculated using Gemma-9B (Team, 2024a) as an oracle model which has more parameters than Llama3-8B. The baseline methods include MPAC (Yoo et al., 2023b), Soft Red List (Kirchenbauer et al., 2023), and SynthID (Dathathri et al., 2024). For MPAC and Soft Red List, the proportion $\gamma$ of green lists is $0.5$ for a balance between detectability and text quality, and also fair comparison with BiMark’s vocabulary bipartiiton setting. For SynthID, the number of tournaments is $30$, as recommended in the default setting. For BiMark, the base scaling factor $\tilde{\delta}$ is $1.0$, and the number of layers $d$ is $10$, which provides a balance between detectability and computational efficiency.
@@
Multi-bit watermarking scenario.
In this experiment, we compare BiMark with MPAC (Yoo et al., 2023b), a state-of-the-art multi-bit watermarking method that is model-agnostic and message-agnostic detectable. We test with varying message lengths (8, 16, and 32 bits) and text lengths (50, 100, 200, and 300 tokens), using message extraction rate, i.e. the bit accuracy between embedded messages and extracted messages, as the evaluation metric. As shown in Tab. 2, MPAC significantly sacrifices text quality when the value $\delta$ added to logit scores of green tokens grows from $1$ to $1.5$, even though it improves the message extraction. In contrast, BiMark consistently achieves higher extraction rates with lower perplexity. For 50-token texts, BiMark improves message extraction rates by 20.87%, 29.50%, and 30.02% for 8-bit, 16-bit, and 32-bit messages respectively, compared to MPAC. The performance improvement is more evident with both shorter texts and longer messages, which typically present greater challenges in watermarking.
We further evaluate resilience against synonym substitution attacks (Jovanović et al., 2023; Zhang et al., 2024d; Hou et al., 2023; Ren et al., 2023) using BERT (Devlin et al., 2018) and WordNet (Miller, 1995). Tab. 3 demonstrates BiMark’s superior performance—for 100-token texts, BiMark maintains 18.44%, 21.30%, and 26.24% higher extraction rates than those of MPAC under 0.1, 0.2, and 0.3 text substitution ratios, respectively. This enhanced resilience is attributed to the fact that the multilayer mechanism provides abundant watermark evidences, leading the statistical patterns of detection to be more distinguishable between watermarked and non-watermarked text.
@@ 5.2 Text Quality Preservation
Watermarks’ impact on generated text quality is assessed in two downstream tasks of language models: text summarization and machine translation. For text summarization, BART-large (Lewis et al., 2019) is employed on the CNN/DailyMail dataset (See et al., 2017), where the performance is evaluated using BERTScore-F1 (Zhang* et al., 2020) and ROUGE-1 (Lin, 2004). For machine translation, MBart (Lewis et al., 2019) is employed on the WMT’16 En-Ro subset (Bojar et al., 2016), where the performance is evaluated using BLEU (Papineni et al., 2002).
Baseline methods include Soft Red List, Gumbel Sampling (Kuditipudi et al., 2023), $\gamma$-Reweight (Hu et al., 2023), DiPmark (Wu et al., 2023), and SynthID (Dathathri et al., 2024). Soft Red List, SynthID, and BiMark use the same settings as previous experiments, while other methods take default or recommended settings in the original works. In Tab. 4, Soft Red List shows obvious quality degradation, where performance drops significantly as watermark strength increases (indicated by $\delta$). Among unbiased methods, BiMark achieves comparable performance while additionally supporting embedding and extracting useful messages. The consistent performance across both summarization and translation tasks demonstrates that our unbiased multilayer watermarking successfully preserves language models’ essential abilities for downstream tasks.
@@
No Watermark 32.45$\pm$ .01 38.32$\pm$ .02 56.21$\pm$.03 21.93$\pm$.17
Soft Red List (1.0) 32.11$\pm$.02 37.97$\pm$.04 55.74$\pm$.18 21.36$\pm$.16
Soft Red List (1.5) 31.61$\pm$.04 37.51$\pm$.06 55.06$\pm$.15 20.67$\pm$.25
Soft Red List (2.0) 31.15$\pm$.02 36.99$\pm$.05 54.17$\pm$.18 19.63$\pm$.02
Gumbel Sampling 32.22$\pm$.03 38.19$\pm$.02 56.12$\pm$.06 22.14$\pm$.05
$\gamma$-Reweight 32.25$\pm$.02 38.09$\pm$ .02 55.67$\pm$.05 21.49$\pm$.14
DiPmark 32.33$\pm$.02 38.21$\pm$.03 56.11$\pm$.04 21.87$\pm$.06
SynthID 32.45$\pm$.03 38.32$\pm$.04 56.17$\pm$.09 22.11$\pm$.03
BiMark (Ours) 32.48$\pm$.03 38.32$\pm$.03 56.14$\pm$.07 22.15$\pm$.09
@@ A.2 Type-II Error Analysis of Single-Layer Unbiased Reweighting
$E[G_{t}]$ =$=\begin{cases}0.5+\tilde{\delta}\tau_{t}&\mathrm{if~{}}0\leq\tau_{t}\leq\frac{1}{1+\tilde{\delta}},\\ 1.5-\tau_{t}&\mathrm{if~{}}\frac{1}{1+\tilde{\delta}}<\tau_{t}<1,\end{cases}$
$\mathrm{Var}[G_{t}]$ =$=\begin{cases}0.25-\tilde{\delta}^{2}\tau_{t}^{2}&\mathrm{if~{}}0\leq\tau_{t}\leq\frac{1}{1+\tilde{\delta}},\\ -0.75+2\tau_{t}-\tau_{t}^{2}&\mathrm{if~{}}\frac{1}{1+\tilde{\delta}}<\tau_{t}<1,\end{cases}$
@@
$G_{t}=\begin{cases}1&\text{if }(e=0\wedge x\in\mathcal{V}_{0})\text{ or }(e=1\wedge x\in\mathcal{V}_{1}),\\ 0&\text{otherwise}.\end{cases}$ (4)
@@
$P_{M,w}(x|x\in\mathcal{V}_{1})$ =$=\begin{cases}(1+\delta)P_{M}(x)&\text{if }e=1\\ (1-\delta)P_{M}(x)&\text{if }e=0\end{cases}$ (5)
$P_{M,w}(x|x\in\mathcal{V}_{0})$ =$=\begin{cases}(1-\frac{\delta\tau}{1-\tau})P_{M}(x)&\text{if }e=1\\ (1+\frac{\delta\tau}{1-\tau})P_{M}(x)&\text{if }e=0.\end{cases}$
@@
$=\sum_{x\in\mathcal{V}}[P_{M,w}^{t}(x|x\in\mathcal{V}_{1},e=1)\Pr(e=1)+P_{M,w}^{t}(x|x\in\mathcal{V}_{1},e=0)\Pr(e=0)$
$\quad+P_{M,w}^{t}(x|x\in\mathcal{V}_{0},e=1)\Pr(e=1)+P_{M,w}^{t}(x|x\in\mathcal{V}_{0},e=0)\Pr(e=0)]G_{t}$
@@
$=0.5\sum_{x\in\mathcal{V}}[P_{M,w}^{t}(x|x\in\mathcal{V}_{1},e=1)+P_{M,w}^{t}(x|x\in\mathcal{V}_{0},e=0)]\qquad\qquad~{}(\because e\sim\text{Bernoulli}(0.5))$
@@
$E[G_{t}]=\begin{cases}0.5+\tilde{\delta}\tau_{t}&\text{if }0\leq\tau_{t}<\frac{1}{1+\tilde{\delta}},\\ 1.5-\tau_{t}&\text{if }\tau_{t}\geq\frac{1}{1+\tilde{\delta}}.\end{cases}$ (6)
To derive $Var[G_{t}]$, since $G_{t}$ takes values of either $1$ or 00, we have $G_{t}^{2}=G_{t}$, and consequently $E[G_{t}^{2}]=E[G_{t}]$. Using the variance formula and this property, we have:
@@
$\mathrm{Var}[G_{t}]:=\begin{cases}0.25-\tilde{\delta}^{2}\tau_{t}^{2}&\text{if }0\leq\tau_{t}<\frac{1}{1+\tilde{\delta}},\\ -0.75+2\tau_{t}-\tau_{t}^{2}&\text{if }\tau_{t}\geq\frac{1}{1+\tilde{\delta}}.\end{cases}$ (7)
@@
Key insights. The Type-II error rate decreases when $E[G_{t}]$ increases or $\mathrm{Var}[G_{t}]$ decreases. When $\tilde{\delta}=1$, $E[G_{t}]$ reaches its maximum when $\tau=0.5$, implying that our watermarking scheme is most effective for high-entropy text, which aligns with previous findings from (Kirchenbauer et al., 2023) and (Kuditipudi et al., 2023). Note that $E[G_{t}]$ reaches its maximum when $\tau=\frac{1}{1+\tilde{\delta}}$. This actually corresponds to the case when $\tau$ is amplified to occupy the entire probability space. In this case, probabilities in $\mathcal{V}_{0}$ will be shrunk to 00. As multilayer reweighting is iterative, this process will concentrate probabilities among a few tokens. This property indicates that shallow layers of reweighting have more impact on the ultimate reweighted probability distribution. Once tokens are shrunk to near-zero probabilities in early layers, even if they are allocated to green lists in deeper layers, their probabilities cannot be effectively recovered.
@@ A.4 Property of XOR operaion
We need to prove that $\Pr(e=1)=1/2$ and $\Pr(e=0)=1/2$, regardless of the value of $x$.
when $x=0$, $e=0\oplus b$. If $b=1$, then $e=0$. If $b=0$, then $e=1$. Because $\Pr(b=1)=\Pr(b=0)=1/2$, we have $\Pr(e=0|x=0)=1/2$ and $\Pr(e=1|x=0)=1/2$.
When $x=1$, $e=1\oplus b$. If $b=1$, then $e=0$. If $b=0$, then $e=1$. Because $\Pr(b=1)=\Pr(b=0)=1/2$, we have $\Pr(e=0|x=1)=1/2$ and $\Pr(e=1|x=1)=1/2$.
@@ Appendix C Algorithms
     $[(\mathcal{V}_{0}^{1},\mathcal{V}_{1}^{1}),(\mathcal{V}_{0}^{2},\mathcal{V}_{1}^{2}),...,(\mathcal{V}_{0}^{d},\mathcal{V}_{1}^{d})]$
@@ 2. If the current previous tokens $\bm{x}_{-h:}$ in the sliding
          4. Let $\theta_{i}=(\mathcal{V}_{0}^{i},\mathcal{V}_{1}^{i},\tilde{\delta})$, and calculate
@@ 5. Obtain the $i$-layer reweighted distribution
     $[(\mathcal{V}_{0}^{1},\mathcal{V}_{1}^{1}),(\mathcal{V}_{0}^{2},\mathcal{V}_{1}^{2}),...,(\mathcal{V}_{0}^{d},\mathcal{V}_{1}^{d})]$
@@
  1. Initialize a ×$\ell\times 2$ voting matrix $\boldsymbol{M}$
@@ Appendix D Examples of Watermarked Text
“We think that it’ll be very helpful in understanding if we’re really getting the benefits,” he said.

### v2

@@ 5.1 Message Embedding Capacity
Experimental setup. For experiments of message embedding capacity, the Llama3-8B model (AI@Meta, 2024) is used for text generation with temperature $1.0$ and top-$50$ sampling. C4-RealNewslike (Raffel et al., 2020) dataset is used as prompts. For pseudorandom operations of watermarking methods, a 2-token sliding context window is used for seeding. For evaluating the generation quality of language models with watermarks, perplexity of generated text is calculated using Gemma-9B (Team, 2024a) as an oracle model which has more parameters than Llama3-8B. The baseline methods include MPAC (Yoo et al., 2024), Soft Red List (Kirchenbauer et al., 2023), and SynthID (Dathathri et al., 2024). For MPAC and Soft Red List, the proportion $\gamma$ of green lists is $0.5$ for a balance between detectability and text quality, and also fair comparison with BiMark’s vocabulary bipartiiton setting. For SynthID, the number of tournaments is $30$, as recommended in the default setting. For BiMark, the base scaling factor $\tilde{\delta}$ is $1.0$, and the number of layers $d$ is $10$, which provides a balance between detectability and computational efficiency.
@@
Multi-bit watermarking scenario. In this experiment, we compare BiMark with MPAC (Yoo et al., 2024), a state-of-the-art multi-bit watermarking method that is model-agnostic and message-agnostic detectable. We test with varying message lengths (8, 16, and 32 bits) and text lengths (50, 100, 200, and 300 tokens), using message extraction rate, i.e. the bit accuracy between embedded messages and extracted messages, as the evaluation metric. As shown in Tab. 2, MPAC significantly sacrifices text quality when the value $\delta$ added to logit scores of green tokens grows from $1$ to $1.5$, even though it improves the message extraction. In contrast, BiMark consistently achieves higher extraction rates with lower perplexity. For 50-token texts, BiMark improves message extraction rates by 20.87%, 29.50%, and 30.02% for 8-bit, 16-bit, and 32-bit messages respectively, compared to MPAC. The performance improvement is more evident with both shorter texts and longer messages, which typically present greater challenges in watermarking.
We further evaluate resilience against synonym substitution attacks (Jovanović et al., 2024; Zhang et al., 2024d; Hou et al., 2024; Ren et al., 2024) using BERT (Devlin et al., 2019) and WordNet (Miller, 1995). Tab. 3 demonstrates BiMark’s superior performance—for 100-token texts, BiMark maintains 18.44%, 21.30%, and 26.24% higher extraction rates than those of MPAC under 0.1, 0.2, and 0.3 text substitution ratios, respectively. This enhanced resilience is attributed to the fact that the multilayer mechanism provides abundant watermark evidences, leading the statistical patterns of detection to be more distinguishable between watermarked and non-watermarked text.
@@ 5.2 Text Quality Preservation
Watermarks’ impact on generated text quality is assessed in two downstream tasks of language models: text summarization and machine translation. For text summarization, BART-large (Lewis et al., 2020) is employed on the CNN/DailyMail dataset (See et al., 2017), where the performance is evaluated using BERTScore-F1 (Zhang* et al., 2020) and ROUGE-1 (Lin, 2004). For machine translation, MBart (Lewis et al., 2020) is employed on the WMT’16 En-Ro subset (Bojar et al., 2016), where the performance is evaluated using BLEU (Papineni et al., 2002).
Baseline methods include Soft Red List, Gumbel Sampling (Kuditipudi et al., 2024), $\gamma$-Reweight (Hu et al., 2024), DiPmark (Wu et al., 2024), and SynthID (Dathathri et al., 2024). Soft Red List, SynthID, and BiMark use the same settings as previous experiments, while other methods take default or recommended settings in the original works. In Tab. 4, Soft Red List shows obvious quality degradation, where performance drops significantly as watermark strength increases (indicated by $\delta$). Among unbiased methods, BiMark achieves comparable performance while additionally supporting embedding and extracting useful messages. The consistent performance across both summarization and translation tasks demonstrates that our unbiased multilayer watermarking successfully preserves language models’ essential abilities for downstream tasks.
@@
[…]
@@ A.2 Type-II Error Analysis of Single-Layer Unbiased Reweighting
[…]
@@
[…]
@@
[…]
@@
$=\sum_{x\in\mathcal{V}}[P_{M,w}^{t}(x|x\in\mathcal{V}_{1},e=1)P(e=1)+P_{M,w}^{t}(x|x\in\mathcal{V}_{1},e=0)P(e=0)$
$\quad+P_{M,w}^{t}(x|x\in\mathcal{V}_{0},e=1)P(e=1)+P_{M,w}^{t}(x|x\in\mathcal{V}_{0},e=0)P(e=0)]G_{t}$
@@
[…]
To derive $\mathrm{Var}[G_{t}]$, since $G_{t}$ takes values of either $1$ or 0, we have $G_{t}^{2}=G_{t}$, and consequently $E[G_{t}^{2}]=E[G_{t}]$. Using the variance formula and this property, we have:
@@
[…]
@@
Key insights. The Type-II error rate decreases when $E[G_{t}]$ increases or $\mathrm{Var}[G_{t}]$ decreases. When $\tilde{\delta}=1$, $E[G_{t}]$ reaches its maximum when $\tau=0.5$, implying that our watermarking scheme is most effective for high-entropy text, which aligns with previous findings from (Kirchenbauer et al., 2023) and (Kuditipudi et al., 2024). Note that $E[G_{t}]$ reaches its maximum when $\tau=\frac{1}{1+\tilde{\delta}}$. This actually corresponds to the case when $\tau$ is amplified to occupy the entire probability space. In this case, probabilities in $\mathcal{V}_{0}$ will be shrunk to 0. As multilayer reweighting is iterative, this process will concentrate probabilities among a few tokens. This property indicates that shallow layers of reweighting have more impact on the ultimate reweighted probability distribution. Once tokens are shrunk to near-zero probabilities in early layers, even if they are allocated to green lists in deeper layers, their probabilities cannot be effectively recovered.
@@ A.4 Property of XOR operaion
We need to prove that $P(e=1)=1/2$ and $P(e=0)=1/2$, regardless of the value of $x$.
when $x=0$, $e=0\oplus b$. If $b=1$, then $e=0$. If $b=0$, then $e=1$. Because $P(b=1)=P(b=0)=1/2$, we have $P(e=0|x=0)=1/2$ and $P(e=1|x=0)=1/2$.
When $x=1$, $e=1\oplus b$. If $b=1$, then $e=0$. If $b=0$, then $e=1$. Because $P(b=1)=P(b=0)=1/2$, we have $P(e=0|x=1)=1/2$ and $P(e=1|x=1)=1/2$.
@@ Appendix C Algorithms
    $(\mathcal{V}_{0}^{1},\mathcal{V}_{1}^{1}),(\mathcal{V}_{0}^{2},\mathcal{V}_{1}^{2}),...,(\mathcal{V}_{0}^{d},\mathcal{V}_{1}^{d})]$
@@ 2. If the current previous tokens $\bm{x}_{-h:}$ in the sliding
[…]
@@ 5. Obtain the $i$-layer reweighted distribution
    $(\mathcal{V}_{0}^{1},\mathcal{V}_{1}^{1}),(\mathcal{V}_{0}^{2},\mathcal{V}_{1}^{2}),...,(\mathcal{V}_{0}^{d},\mathcal{V}_{1}^{d})]$
@@
 1. Initialize a $\ell\times 2$ voting matrix $\boldsymbol{M}$
@@ Appendix D Examples of Watermarked Text
“We think that it’$l$ be very helpful in understanding if we’re really getting the benefits,” he said.